With celebrity​ bonds, celebrities raise money by issuing bonds to investors. The royalties from sales of the music are used to pay interest and principal on the bonds. In April of​ 2009, EMI announced that it intended to securitize its back catalogue with the help of the Bank of Scotland. The bond was issued with a coupon rate of 6.9​% and will mature on this day 36 years from now. The yield on the bond issue is currently 6.1​%. At what price should this bond trade​ today, assuming a face value of ​$1,000 and annual​ coupons?
Answer #1

Face Value = $1000

Annual Coupon payment = $1000*6.9%

= $69

No of years to maturity from today(n) = 36 years

Yield to maturity (YTM) = 6.1%

Calculating the Price of Bond:-

+ ... Coupon Payment (1+YTM) Price Coupon Payment Coupon Payment + (1+YTM) (1 + YTM2 FaceValue (1+YTM)

Price = \frac{69}{(1+0.061)^{1}}+\frac{69}{(1+0.061)^{2}}+.....+\frac{69}{(1+0.061)^{36}}+\frac{1000}{(1+0.061)^{36}}

Price = $996.9465 + $118.644

Price = $1115.59

So, Price the bond should trade at $1115.59

If you need any clarification, you can ask in comments.     

If you like my answer, then please up-vote as it will be motivating